A rope is tied to the top of a 10 m flag pole. The rope is tensioned to 120N and then tied to a stake in the ground 20m from the bottom of the pole. The ground is level. Find the vertical and horizontal components of this force.


1
Expert's answer
2021-01-16T17:23:49-0500


The situation is shown in the figure. From the right triangle "\triangle ABC", obtain:

"AB = \sqrt{CB^2 + AC^2} = \sqrt{10^2 + 20^2} = 10\sqrt{5}m"

The vertical projection of the force "\mathbf{F}" is:


"F_v = F\sin\theta"

According to the sine definition, obtain:


"F_v = F\dfrac{CB}{AB} = 120\dfrac{10}{10\sqrt{5}} = \dfrac{120}{\sqrt{5}} = 24\sqrt{5}N"

The horizontal projection is (according to the cosine definition):


"F_h = F\cos\theta= F\dfrac{AC}{AB} = 120\dfrac{20}{10\sqrt{5}} = 48\sqrt{5}N"

Answer. Horizontal "48\sqrt{5}N", vertical "24\sqrt{5}N".